40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Thu Mar 4 13:37:14 2021, aaron, DailyProgress, Electronics, PSOMA S mA->MHz, ITC 502 current noise mA_to_MHz.pdf002F0263-0615-4566-9192-F57A14C8B543.jpegITC502_ASD.pdf
    Reply  Sun Mar 7 19:32:32 2021, rana, DailyProgress, Electronics, PSOMA S mA->MHz, ITC 502 current noise 
    Reply  Wed Mar 10 16:03:19 2021, aaron, DailyProgress, Electronics, current noise 
       Reply  Thu Mar 11 13:56:47 2021, aaron, DailyProgress, Electronics, current noise wire_shelf_5.jpg
Message ID: 2673     Entry time: Thu Mar 4 13:37:14 2021     Reply to this: 2675   2681
Author: aaron 
Type: DailyProgress 
Category: Electronics 
Subject: PSOMA S mA->MHz, ITC 502 current noise 

enter Thu Mar 4 13:37:09 2021

Picked up from Downs (see PSOMA hardware inventory for specifications):

  • label maker and tape
  • 4x longpass filters

current to frequency

  1. Plug in windows laptop to Teraxion driver box
  2. Turn on driver box, then 'connect' to driver from the GUI. The interlock automatically switches to 'on' and the laser powers up at 1550.12 nm. I didn't change any settings on the Teraxion laser.
  3. Power from the blue fiber carrying Teraxion beam to PSOMA table puts out 0.6 mW. I replace the PSOMA N beam with the Teraxion beam just before the Faraday isolator.
  4. With the PSOMA South laser current drive at 149.36 mA and TEC at 9.305 kOHm, the beat is at 2.078 GHz (HP 8560E). I tune the S laser to 120.74 mA to get the beat to around 50 MHz.
    • There seems to be some thermalization happening with either or both lasers, because the beat note is moving monotonically at about 200 kHz/s. I'm searching for an SR554 in the meantime.
    • Didn't locate a transformer amplifier. I took the following data by hand, but as you can see by the apparent hysterisis, the laser frequency is still monotonically changing.
    • Second set of data after the beat note settled down
current (mA) 118.30 118.11 117.68 117.56 117.35 117.50 117.59 117.65 117.74 117.83 118.04 118.11 118.29 118.47 118.55 118.62 118.76 118.89
beat frequency (MHz) 148 120 58 42 9 35 47 55 69 83 114 123 149 175 186 197 2162 235
current (mA) 117.89 117.73 117.56 117.41 117.22 117.13 117.07 117.01 117.13 117.23 117.35 117.63 117.76 117.95 118.06 118.11 118.20 118.26
frequency (MHz) 137 112 88 66 38 26 17 7 26 40 58 101 120 148 164 170 183 192

ITC 502 current noise

  1.  Though I don't have a transformer, I expect about 1 nA/rtHz from the ITC502 I'm measuring. Driving a 20 Ohm dummy load with a , this should be 20 nV/rtHz at the input of the SR560 -- well above its 4 nV/rtHz noise floor. Seems fine for now.
    • Repeating the measurement here, with the following pin mapping
      • 1, 5 have a dummy load to lift the interlock
      • 3, 7 are sending current across a 20 Ohm dummy load
      • SR560 is on battery power measuring in AC-coupled, low noise mode with G=100 and a 100 kHz low pass filter (in addition to the implicit 0.03 Hz high pass due to AC coupling).
      • Moku is set to 'data acquisition' mode, and the signal is sent to DC-coupled 50 Ohms with 1 Vpp range. Acquiring 7 min data at 256 kS/s.
    • After setting up the moku for remote access (below), I killed the lights, left the lab, and ran the following from the directory where I wanted the data file stored:
      • python [path_to_script]mokuDataLogger.py -c ch2 -i 10.0.5.220 -d 420 -s 256e3 --useInternal --fileType bin

Update: Plotted the current noise ASD, using the 'modified welch' method found In labutils/moku/modifiedPSD. I used 'median' averaging for the welch spectrograms, and the script from Anchal and Scott improves averaging at high frequencies by building the spectrum frequency decade by deacde, and calling welch with as much averaging as possible for the chosen decade. Shaded region is the 1 sigma confidence around the median. The right axis shows the current noise referred to the S laser's frequency noise, assuming 147.86 MHz/mA.

Hm, not sure if it makes sense yet.

Update: No, it did not make sense. I converted A^2/Hz to A/rtHz as if my data were already in dB, but they were in magnitude units; and the twinned y axis for frequency noise was not log scaled. I've updated the plot attached and the script in the git repo, and added a line where the noise floor from voltage noise of SR560 across 20 Ohms should be (not measured). I'm still surprised to see current noise so far above the nominal noise floor of the SR560.

miscellany

  • Opened one of the SR560 lsthat was constantly hitting overload in either DC or GND mode.
    • front panel offset was turned too far. Adjusted offset as suggested in calibration guide until pin 14 on U407 was at the same voltage as unit's ground (was ~6V).
    • Unit still overloading, identified that one leg of U106 (front end op amp) was not at 0V. Manual suggests this constant overloading is a sign U106 is busted.
  • uninstalled pip and conda on spirou, which were improperly installed by sudo user. All python on spirou now through controls user installed anaconda. Also initialized moku and psoma environments with some of the usual packages. Pip is installed in the virtual envs (not in base), and all pip installations should occur after conda package installations.
    • sudo python -m pip uninstall pip
      conda activate moku
      conda install pip
      pip install pymoku
      moku update fetch
    • sudo apt-get install libavahi-compat-libdnssd1
  • connected moku to ethernet on router port 10.0.5.220
    • first need to configure moku to accept ethernet connection
      • when connecting via ipad, click gear icon before 'go'ing to splash page listing available instruments.
      • Turn 'on' connection via ethernet. I used DHCP, and was assigned IP 192.168.73.1

Returned ITC502 to driving S laser, still below 120 mA.

exit Thu Mar 4 23:04:28 2021

Attachment 1: mA_to_MHz.pdf  28 kB  Uploaded Thu Mar 4 22:45:56 2021  | Hide | Hide all
mA_to_MHz.pdf
Attachment 2: 002F0263-0615-4566-9192-F57A14C8B543.jpeg  318 kB  Uploaded Thu Mar 4 23:51:52 2021  | Hide | Hide all
002F0263-0615-4566-9192-F57A14C8B543.jpeg
Attachment 3: ITC502_ASD.pdf  47 kB  Uploaded Mon Mar 8 19:03:09 2021  | Hide | Hide all
ITC502_ASD.pdf
ELOG V3.1.3-